### v1

# Anomalous Ionization in the Central Molecular Zone by sub-GeV Dark Matter

Abstract
We demonstrate that the anomalous ionization rate observed in the Central Molecular Zone can be attributed to MeV dark matter annihilations into $e^{+}e^{-}$ pairs for galactic dark matter profiles with slopes $\gamma>1$. The low annihilation cross-sections required avoid cosmological constraints and imply no detectable inverse Compton, bremsstrahlung or synchrotron emissions in radio, X and gamma rays. The possible connection to the source of the unexplained $511$ keV line emission in the Galactic Center suggests that both observations could be correlated and have a common origin.
I introduction
Various measurements are known to indicate an unexpectedly high rate of H2 ionization rates around the Galactic Center (GC), particularly in the Central Molecular Zone (CMZ). This has been evidenced by a variety of tracers, including methanol production Coutens et al. (2017), (indicating an ionization rate of $\zeta\sim 10^{-15}$ s-1), and the observation of $\textrm{H}_{3}^{+}$ lines Oka and Geballe (2020); Indriolo et al. (2009); Goto et al. (2014) (requiring an ionization rate around $\zeta\sim 2\times 10^{-14}$ s-1Le Petit et al. (2016); Ravikularaman et al. (2024); Geballe and Oka (2010) and other ionized molecules, such as $\textrm{H}_{3}$+ Indriolo et al. (2015); van der Tak et al. (2006) (indicative of an ionization rate $\gtrsim 10^{-15}$ s-1).
Recently, Ref. Ravikularaman et al. (2024) showed that the diffuse flux of cosmic rays (CRs) falls short of explaining such high ionization rates in the CMZ, arguing that it requires an extremely large injection power for CRs at the GC, of around $10^{40}-10^{41}$ erg s-1 (see also Ref. Phan et al. (2018)). Moreover, it was shown that such a CR flux would exceed the MeV gamma-ray emission observed in the inner galaxy ($|l|<$ $30^{\circ}$ and $|b|<$ $15^{\circ}$). This may indicate the presence of a source highly concentrated around the GC that is producing low energy CRs that lose their energy very quickly, via ionization of molecular gas Indriolo et al. (2009). From the latter, one can infer that these particles must be injected at low energies, to efficiently be confined around the CMZ thereby preventing them from propagating long distances.
Here, we investigate whether annihilation of sub-GeV DM Boudaud et al. (2017); Cirelli et al. (2023a) into $e^{\pm}$ pairs can explain the observed ionization rate for DM masses below $\sim 100$ MeV, without being in conflict with other known current constraints. While DM density distributions with a central spike could also provide promising results, in this work we consider several dark matter profiles including the Navarro-Frenk-White (NFW) Cirelli et al. (2024) and those with larger slope $\gamma$.
Figure 1: Maximum ionization rate that can be provided by sub-GeV DM, allowed by CMB constraints on $\langle\sigma v\rangle$ Lopez-Honorez et al. (2013); Slatyer (2016). We compare the predicted ionization rate obtained with an analytical calculation and with a numerical calculation using a dedicated CR propagation code (DRAGON2). The shaded area represents the ionization rate observed in the CMZ.
The motivation for sub-GeV DM is strong, and has previously been discussed in detail in the literature, see e.g. Refs. Battaglieri et al. (2017); Cirelli et al. (2021); Balan et al. (2024a) and Ref. Cirelli et al. (2024) for a recent review. Presently, CMB constraints Lopez-Honorez et al. (2013); Slatyer (2016), provide the most robust constraints for annihilation of sub-GeV DM particles, and, as shown in Fig. 1, these are consistent with a DM origin of the anomalous ionization rate observed in the CMZ for profiles generally cuspier than a NFW DM profile Navarro et al. (1996) and low DM masses (and even allow for combinations of mass and DM profiles that could exceed the observed ionization rate), due to the fact that the (low-energy) DM products can deposit roughly all their energy within the CMZ. Here we investigate whether Galactic probes are able to constrain and rule out this explanation of the CMZ ionization rate.
Among the most interesting probes for sub-GeV DM is the $511$ keV line emission, that forms from the decay of positronium formed by the positrons in the Galaxy, and was shown to produce strong constraints on sub-GeV DM annihilation into Standard Model states De la Torre Luque et al. (2023a); Vincent et al. (2012). In fact, $511$ keV line emission from the bulge of the Galaxy remains a mystery Prantzos et al. (2011); Kierans et al. (2019); Siegert (2023) and sub-GeV DM was proposed as a possible explanation Boehm et al. (2004). The observation of the line at the bulge requires a steady production of positrons at a rate of a few times $10^{50}$ $e^{+}$/yr Siegert et al. (2016); Kierans et al. (2020). Once the electrons and positrons are produced, they lose energy, mainly via ionization of the molecular gas in the CMZ. If the medium is dense enough, as is the case for the CMZ, they release all their energy into the medium after travelling less than $\sim 100$ pc (for $e^{\pm}$ with energies below a few tens of MeV for standard diffusion parameters found in CR analyses). Once thermalized, the relativistic positrons produce positronium bound states, in the form of a singlet state, called para-positronium (p-ps), that decays into two gamma rays each with $511$ keV energy, or in a triplet state, called ortho-positronium (o-ps), that decays in $3$ photons Badertscher et al. (2007) producing continuum emission below $511$ keV. We will also discuss the possibility that both observations are connected and explore whether the parameter space of sub-GeV DM that can explain both effects is currently ruled out.
There is another long-standing anomalous emission from the GC that has been related to DM Goodenough and Hooper (2009); Hooper and Goodenough (2011) and other leptonic sources – from milisecond pulsars (MSPs) Abazajian and Kaplinghat (2012); Yuan and Zhang (2014) to leptonic outbursts Cholis et al. (2015): The GeV Fermi-LAT Galactic Center Excess (GCE) Ackermann et al. (2017). This excess spatially coexists with the CMZ (although it is more extended) and its distribution can be well fitted with a contracted NFW DM profile Di Mauro (2021) (c-NFW). The analyses of Fermi-LAT data suggest a luminosity of $10^{37}$ erg s-1 to explain the GCE. As we discuss in this work, it is quite unlikely that the source of this excess can be linked to the other anomalies described above, but understanding these anomalous observations can allow us to understand better the dynamics of the GC.
The remainder of this paper is as follows: We first detail our estimates of the CMZ ionization rate from the annihilation of sub-GeV DM particles. Then, we test if the observations of the CMZ ionization rate are compatible with cosmological constraints and if there would be any associated significant inverse Compton or bremsstrahlung emission. Finally, we discuss whether the CMZ ionization rate and the $511$ keV line emission, could share a common (DM annihilation) origin and comment on the potential of finding correlations between these two observables.
II Ionization rate from sub-GeV DM annihilation in the CMZ
The CMZ, located around the GC, is an approximately cylindrical region around $200$ pc wide and $100$ pc high Ferriere et al. (2007); Ravikularaman et al. (2024), of high-density molecular gas ($n_{\textrm{H}}\sim 150$ cm-3). Low energy CRs $\lesssim 100$ MeV are subject to severe ionization energy losses that limit their propagation (see Fig. 3). At higher energies, these particles can escape the CMZ and produce potentially detectable secondary diffuse emission (such as bremsstrahlung, inverse Compton or synchrotron radiation).
The ionization rate $\zeta$ caused by the electrons and positrons of kinetic energy $E$ at position vector x inside the CMZ can be calculated using Eq. (1) below Padovani et al. (2009); Phan et al. (2023).
[…]
where we the factor $2$ accounts for both $e^{-}$ and $e^{+}$, $E_{\textrm{max}}$ is the maximum kinetic energy with which the electrons or positrons are injected and $E_{\textrm{min}}$ is the minimum kinetic energy to produce ionization in the gas ($13.6$ eV), $\sigma(E)$ is the $e$-H2 cross-section, calculated as in Eq. (7) of Ref. Padovani et al. (2009), $\theta_{e}(E)$ is the number of secondary ionizations per primary ionization, calculated as Eq. (2.23) of Ref. Krause et al. (2015), and $J(E)$ is the flux that can be computed as
$J(E,\textbf{x})=\frac{d\phi_{e}}{dE}\left(E,\textbf{x}\right)\frac{\beta_{e}(E)c}{4\pi}\,,$ (2)
where $\beta(E)c$ is the speed of the electron or positron and $\frac{d\phi_{e}}{dE}$ is their differential flux in the CMZ. We compute the differential flux of electrons and positrons induced by sub-GeV DM annihilation in two ways: an analytical approximation that allows us to estimate the diffuse $e^{\pm}$ flux within the CMZ, and a numerical calculation that allows us to compute their all-sky diffuse emission. We also note that these annihilations are expected to also produce ionizing photons through final state radiation, however, the cross-section for ionizing H2 is $\sim 10^{-6}$ times lower than from electrons/positrons at $1$ MeV, and negligible in comparison with the ionization caused by $e^{\pm}$.
Analytical approximation — In this approximation, we neglect diffusion of the positrons and electrons within the CMZ and calculate their steady-state distribution as
[…]
with $b_{e}$ as the energy loss term ($b_{e}\equiv\frac{dE}{dt}^{\textrm{ion}}$), calculated as implemented in the DRAGON2 code (see Eq. (C.39) of Ref. Evoli et al. (2017)), $n(x_{\textrm{CMZ}})=150$ cm-3 is taken as the mean density of the CMZ Ravikularaman et al. (2024) and $Q_{e}\left(E,\textbf{x}\right)$ as the injection spectrum (source term) from DM annihilation, which takes the following form Cirelli et al. (2011)
$Q_{e}\left(E,\textbf{x}\right)=\frac{\langle\sigma v\rangle}{2}\left(\frac{\rho_{\chi}(\textbf{x})}{m_{\chi}}\right)^{2}\frac{dN_{e}^{\textrm{ann}}}{dE_{e}}\,,$ (4)
where $\langle\sigma v\rangle$ is the DM thermally averaged annihilation cross section, $m_{\chi}$ is the DM mass, $\rho_{\chi}$ is the DM density in the CMZ and $\frac{dN_{e}^{\textrm{ann}}}{dE_{e}}=\delta(E+m_{e}-m_{\chi})$.
We notice that the approximation used to obtain Eq. (3) is only valid for low energy particles below a few tens of MeV and in very dense media, but the simplifying advantage of this estimation is that it does not depend on the diffusion coefficient or other propagation parameters that are not precisely known near the GC.
Numerical calculation — In our main approach, we compute the steady-state diffuse $e^{\pm}$ spatial distribution and energy spectra with a recent version De La Torre Luque (2023) of the DRAGON2 code Evoli et al. (2017, 2018), a dedicated CR propagation code designed to numerically solve the full diffusion-advection-reacceleration-loss equation for the transport of charged particles in the Galactic environment Ginzburg and Syrovatskii (1969). We use the same prescription and propagation parameters as used in Refs. De la Torre Luque et al. (2023b, c, d, a), where we model the fluxes of sub-GeV positrons from DM and other exotic positron sources using a set of propagation parameters obtained through a combined fit to CR measurements by AMS-02 Aguilar et al. (2021, 2018, 2013) and Voyager-1 Webber et al. (2002); Stone et al. (2013).We refer to Ref. De la Torre Luque et al. (2023d) for full details.
The advantage of this approach is that it is able to account for the effects of diffusion and reacceleration, but it has two main caveats: the first is that it depends on the extrapolation of the propagation parameters obtained from analyses of local CR data at GeV energies (see Ref. De la Torre Luque et al. (2023d) for detailed discussion), and the second is that the code does not consider how the dynamics of diffusion can change when crossing the border of the CMZ. Regarding the latter, Ref. Phan et al. (2023) showed that crossing a very dense medium can lead to a reduction of the flux of particles within molecular clouds, thereby reducing the ionization rate. This effect is more pronounced for lower positron energies and hence lower DM masses since they produce lower energy positrons. However, we note that this effect is not so relevant in our case, given that the electrons and positrons producing ionization come mostly from the interior of the CMZ, and that even with high suppression of the $e^{\pm}$ flux the predicted ionization rate can exceed that observed at the CMZ by orders of magnitude, we will be discussed further below. Therefore, differences in the predicted ionization rates from both approaches are expected, especially at masses above a few tens of MeV where the analytical approximation is not valid anymore because the diffusion rate starts to become dominant (see Fig. 3).
III Results
The main constraints on sub-GeV DM arise from cosmological observations of the early Universe such as the effective number of relativistic species $N_{\textrm{eff}}$ and the requirement for successful Big Bang Nucleosynthesis (BBN). These observations can robustly rule out DM masses below around $1$ MeV Boehm et al. (2012, 2013); Nollett and Steigman (2015), and even slightly above, depending on certain assumptions. However, recent work such as Refs. Escudero (2019); Sabti et al. (2020), has shown compatibility with observation for DM masses above a few MeV, as long as the dominant early Universe channel for DM annihilation is into neutrinos. For higher masses, there are strong astrophysical and cosmological constraints (see a recent review Ref. Balan et al. (2024b) or a more comprehensive revew of DM candidates Ref. Cirelli et al. (2024)). The strongest constraints so far come from XMM-Newton measurements Cirelli et al. (2023b); De la Torre Luque et al. (2023d) of the X-ray Galactic diffuse emission, impact of the $e^{\pm}$ injection from DM annihilation on the CMB anisotropies Slatyer (2016); Lopez-Honorez et al. (2013), and the longitude profile of the $511$ keV line De la Torre Luque et al. (2023a).
The estimated maximum ionization rate from sub-GeV DM with mass and cross-section combinations that are allowed by CMB constraints (taken from the bottom left panel of Fig. 4 of Ref. Slatyer (2016), which are obtained from the impact of DM annihilation on CMB anisotropies) is shown in Fig. 1, where we represent in the $y$-axis the predicted ionization rate from sub-GeV DM in the CMZ for generalized NFW profiles with a slope ranging from $\gamma=1$ (standard NFW distribution), to $\gamma=1.5$ (a Moore profile Moore et al. (1999)), including what is known as a contracted-NFW (cNFW) profile, which corresponds to a $\gamma=1.25$, which is renown for being the best-fit profile for the GeV Fermi-LAT Galactic Center Excess (GCE) Ackermann et al. (2017); Di Mauro (2021). In this figure, the shaded region represents the band of ionization rate values measured at the CMZ. Dashed lines represent the predicted ionization rate for the analytical approximation, while the solid lines represent the result obtained with the numerical calculation using DRAGON2. The latter should be more representative, and we have checked that uncertainties from the diffusion parameters can affect the benchmark prediction by at most a factor of a few (see Ref. De la Torre Luque et al. (2023d)). As a consequence, our result shows that there is a large region of parameter space (of DM profile slopes and DM masses) that can explain the observed ionization rate without violating cosmological constraints.
We notice that both approaches used to compute the ionization rate are approximate, given the difficulties of treating diffusion in the CMZ, however, the parameter space accessible to explain the measured ionization rate is so large that our conclusion should remain robust. To provide some quantitative examples, consider a $\simeq 10$ MeV DM particle annihilating at a rate of $\langle\sigma v\rangle\simeq 5\times 10^{-30}$, this can produce an ionization rate of $\zeta\simeq 2\times 10^{-14}$ s-1 for a slope of $\gamma=1.25$. For the same slope, a $1$ MeV DM particle could produce the observed ionization rate for $\langle\sigma v\rangle\simeq 2\times 10^{-33}$. For a Moore profile ($\gamma=1.5$) this ionization rate is achieved for a $\langle\sigma v\rangle\simeq 1.25\times 10^{-35}$ and $m_{\chi}=10$ MeV. For ease of comparison, we show in Fig. 2 (right axis – dashed lines) the expected ionization rate for a fixed annihilation cross section of $\langle\sigma v\rangle=10^{-35}$ cm3/s, with the numerical diffusion calculation (shown also in a single plot in Fig. S2 of the SM).
In terms of what concerns constraints from Galactic probes, we remark that X-ray constraints are competitive with those from CMB, although for DM particles with masses below $\sim 10$-$20$ MeV, diffuse reacceleration of the $e^{\pm}$ pairs injected by DM De la Torre Luque et al. (2023d); Seo and Ptuskin (1994), induces uncertainties at MeV energies. We have tested the X-ray emission produced by the DM-induced $e^{\pm}$ population using the Hermes code Dundovic et al. (2021) (as in Ref. De la Torre Luque et al. (2023d)), with the same propagation parameters as used for the estimations of the ionization rate in the CMZ, finding that XMM-Newton observations cannot rule out DM with masses below a few tens of MeV, as shown in the left panel of Fig. 5. We have also calculated the bremsstrahlung emission (using the Hermes code as in Ref. Luque et al. (2024)) at the MeV scale. Although bremsstrahlung emission is expected to be more important for low DM masses Cirelli et al. (2013) and dominates over the inverse Compton emission for DM masses below a few tens of MeV Bartels et al. (2017), we find that this emission is not expected to be significant for the annihilation strength needed to explain the CMZ ionization rate and it does not represent an important constraint (see right panel of Fig. 5). As expected, the bremsstrahlung emission is not competitive with observables such as $511$ keV line, X-rays or the CMB.
IV Possible connection with the 511 keV line emission
We now investigate the tantalizing possibility that the $511$ keV line measurements around the GC are correlated with the hydrogen ionization rate in the CMZ. Our discussion is motivated by the following:
i) Refs. Beacom and Yuksel (2006); Sizun et al. (2006, 2007) have found that the source of the bulge $511$ keV emission must require positrons injected with energies lower than a few MeV, so that only DM with mass below $\lesssim 10$ MeV is able to produce the emission.
ii) Ref. De la Torre Luque et al. (2023a) showed that, due to the propagation of positrons, the longitude profile of the $511$ keV line emission can be reproduced only for DM profiles cuspier than NFW, also favouring low DM masses.
iii) The power density from DM annihilations (Eq. 5) in the CMZ is compatible with observations of the line emission at the GC for a cross-section $\langle\sigma v\rangle\lesssim 10^{-32}$ cm3/s. For example, a $1$ MeV particle in a Moore DM profile can produce $\sim 10^{52}\leavevmode\nobreak\ \rm erg/s\sim 5\cdot 10^{50}e^{+}$/yr for $\langle\sigma v\rangle\lesssim 10^{-33}$ cm3/s.
$\frac{dE}{dt}\sim\frac{1}{2}\frac{\rho_{DM}^{2}}{m_{DM}}\langle\sigma v\rangle\,,$ (5)
To test if both can be connected, we compute the $511$ keV emission in the bulge (integrated in the inner $8^{\circ}$) from the annihilation of sub-GeV DM for the different DM profiles considered in this work, following the same approach as in Ref. De la Torre Luque et al. (2024), which is consistent with previous evaluations Calore et al. (2021, 2022); De la Torre Luque et al. (2023b, c); Carenza et al. (2024). Adopting an annihilation rate of $\langle\sigma v\rangle=10^{-35}$ cm3/s, we depict the predicted $511$ keV emission in Fig. 2 (left axis – solid lines), where we include a shaded gray band that accounts for the observed emission at the bulge Siegert et al. (2016), to which we add a conservative $50\%$ uncertainty (to account for underestimated systematic uncertainties). From a comparison to the ionization rate at the CMZ (right axis of the figure – dashed lines), we observe that the annihilation rate needed to explain the $511$ line emission at the bulge is lower than the annihilation rate that explains the CMZ ionization rate, by at least a factor of a few, meaning that a simultaneous explanation of both phenomena seems difficult under the DM hypothesis. This may indicate that the source of the CMZ ionization rate is not symmetrically injecting of electrons and positrons, and could instead be dominated by ionizing protons (or only electrons).
However, we note here that while uncertainties from propagation parameters similarly affect both predictions, the estimated $511$ keV emission is also affected by the conditions for positronium production. The formation of positronium from thermal positrons largely depends on the composition, temperature and density of the medium Guessoum et al. (1991), hindering our ability to make a robust prediction of this quantity (especially around the GC). Therefore, it would be difficult to firmly rule out the DM hypothesis of the CMZ ionization rate from the estimated $511$ keV line bulge emission given that uncertainties in this estimate are larger than an order of magnitude and that the cross-sections needed to match both observations can be close (e.g. for the cNFW profile case with 1 MeV DM where the 511 keV emission is compliant at $\phi^{511\,\textrm{keV}}_{8^{\circ}}\gtrsim 10^{-3}$ cm-2 s-1 and $\zeta\simeq 10^{-16}\,$s-1, only a factor of a few below the observed rate shaded red). In addition, we also remark that there are other mechanisms that could boost the ionization rate without affecting the $511$ keV production: invoking more exotic DM scenarios (that have already been proposed in the literature Farzan and Rajaee (2020); Feng and Kang (2013); Sheng et al. (2024)), it is possible to have asymmetric DM products, i.e. a different amount of electrons and positrons produced. Another intriguing possibility that may make both anomalies compatible with DM is that the direct interaction of DM particles with the H2 molecules also produces ionization.
Figure 2: Left axis: Predicted $511$ keV bulge emission associated with the $e^{\pm}$ population from annihilations for alternative DM density profiles (solid lines). The shaded band shows the constraint set by the observed emission from the bulge, with conservative $50\%$ uncertainties. Right axis: Expected ionization rate in the CMZ from annihilation of sub-GeV DM particles into $e^{+}e^{-}$, as obtained with the numerical calculation (dashed lines). The red band illustrates the range of observed values of ionization rate in the CMZ.
Since correlations between the $511$ keV line emission and the ionization rate in the CMZ can only happen if the source is injecting simultaneously electrons or positrons, we remark that observations of the ionization rate in the CMZ or even in molecular clouds, could help to discern the nature of the source producing the puzzling $511$ keV flux around the GC. However, it is clear that any source producing the emission of $511$ keV photons at the GC must be also play a role in ionizing the CMZ. These correlations will be explored more thoroughly in a separate work.
Finally, one may also wonder if the GCE could explain or significantly contribute to the CMZ ionization rate. Ref. Silk (2018) already argued about the possible connection between the GCE and the ionization rate in the CMZ caused by ultracompact minihalos of self-annihilating DM. However, the current observations point to a low-energy source of electrons or positrons to explain the ionization rate at the CMZ, that could not explain the GCE, that is observed at GeV energies. This holds both if the source is DM annihilating or millisecond pulsars. To demonstrate this, we have computed the ionization rate in the CMZ that is produced by a DM particle that can fit the GCE (to be concrete, we take, from Ref. Calore et al. (2015), the $\sim 10$ GeV DM particle that reproduces the GCE in the $\tau$ channel for a DM distribution with slope of $\gamma=1.2$), which is of $\zeta\sim 10^{21}$ s-1  [^1] We also computed the ionization rate provided by a $\sim 50$ GeV DM particle annihilating into the $b\bar{b}$ channel that fits the GCE Calore et al. (2015) and found that this candidate would produce a lower ionization rate in the CMZ, roughly by a factor of $2$-$3$, which is likely due to the higher DM mass (therefore, lower energy products).. In addition, we have computed the ionization rate produced by a population of millisecond pulsars distributed following a NFW distribution with slope of $\gamma=1.2$, whose emission of $e^{\pm}$ is a power-law with spectral index of $2$, cut-off at $3$ GeV and total power of $10^{37}$ erg/s (that could be adequate to explain the high energy tale of the GCE, as computed in Ref. Petrović et al. (2015)). In this case, the ionization rate is also largely insufficient to explain the measurements of ionization at the CMZ: $\zeta\sim 10^{23}$ s-1. Therefore, it seems that any of the currently proposed sources of the GCE would produce a negligible ionization rate in the CMZ, even lower than that from the diffuse CR sea Phan et al. (2023).
V Discussion and conclusion
In this Letter, we show that the anomalous ionization rate observed at the CMZ can be explained by DM particles of mass below a few tens of MeV annihilating into $e^{+}e^{-}$, for DM distributions with slopes slightly above that of the NFW $\gamma>1$. The thermally-averaged annihilation cross-section required to match the observed ionization rates at the CMZ is below $10^{-33}$ cm3/s for a contracted NFW distribution ($\gamma=1.25$), and can be as low as $10^{-37}$ cm3/s for a Moore DM distribution, all well below the most stringent existing upper limits. Using both analytical approximate calculations and a more detailed numerical calculation with the DRAGON2 code, we show that the required annihilation cross-sections needed to explain the CMZ ionization rate are so low that annihilation of sub-GeV DM would produce no detectable inverse Compton, bremsstrahlung or synchrotron emission. The fact that sub-GeV DM can even exceed the observed ionization rate in the CMZ without being in conflict with these probes demonstrates that this observable can be a very powerful tool for constraining sub-GeV DM. We find that the annihilation cross section needed to explain the CMZ ionization rate is above, by at least a factor of a few, that needed to explain the bulge $511$ keV emission, whilst emphasizing that the uncertainties in the 511 keV emission make it difficult to rule out DM as the source of ionization in the CMZ presently. Therefore, we determine that the $511$ keV emission can not conclusively rule out the DM explanation of the CMZ ionization rate for DM masses below a few MeV, leaving open a window to explore if both anomalies are related (even beyond the DM hypothesis). However future experiments, most notably the imminent launch of NASA’s COSI space telescope in 2027 Siegert et al. (2020), should clarify these issues. We conclude that the annihilation of sub-GeV DM that can explain the ionization rate in the CMZ could significantly contribute to an outstanding astrophysical mystery, namely the nature of the $511$ keV line emission. Finding spatial correlations between molecular clouds and positron line emission, as well as X-ray, diffuse gamma ray and radio constraints on correlated emissions may enable us to demonstrate that the source of molecular ionization must be leptonic, i.e. via injection of electrons and positrons, in deeply embedded regions of the GC.
VI Acknowledgements
We would like to thank Marco Cirelli for helpful feedback on the draft of this letter. SB is supported by the STFC under grant ST/X000753/1. SB is supported by the STFC under grant ST/X000753/1. PDL is supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union ”NextGenerationEU”/PRTR. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, both funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the Swedish National Infrastructure for Computing (SNIC) under project No.2022/3-27 partially funded by the Swedish Research Council through grant no. 2018-05973.
Appendix A Relevant timescales
Our calculations include all the relevant Galactic CR propagation processes in a broad energy range. For signals from sub-GeV DM, it is crucial to account for diffusion, reacceleration, and energy losses, that become dominant below a few tens of MeV. The most relevant cooling processes at these energies are Coulomb and ionization losses, although we include in our simulation synchrotron, bremsstrahlung, inverse Compton and catastrophic losses (annihilation) as well. A comparison of the relevant timescales is illustrated in Fig. 3, where we show a band around the benchmark diffusion time (i.e. diffusion time of particles with our benchmark diffusion setup) to indicate the difference between the optimistic and pessimistic scenarios. Moreover, cooling is much faster than diffusion. We note that the average distance travel by a positron injected at $10$ MeV (which can be calculated as $\langle l_{\text{travel}}\rangle\approx\sqrt{2D\tau}$, being $D$ the diffusion coefficient and $\tau$ the dominant timescale) is larger than $100$ pc in a $1$ cm-3 gas, without accounting for wind advection or reacceleration (that can boost $\langle l_{\text{travel}}\rangle$ to even more than $500$ pc).
Figure 3: Main timescales involved in the propagation of electrons and positrons generated by sub-GeV DM particles, for a gas density of $1$ cm-3 composed by H and a $10\%$ of He. Gas densities. The magnetic field strength here is set to $5$ $\mu$G and the energy density of the interstellar radiation fields is taken from Evoli et al. (2020). The grey band indicates the difference in diffusion time in the Optimistic and Pessimistic setups, used to calculate uncertainties in the DM bounds.
Appendix B Ionization rate for fixed cross sections
For ease of comparison, we show in Fig. 4 the predicted ionization rate in the CMZ from sub-GeV DM annihilation, for different DM distributions and a fixed annhilation cross-section of $\langle\sigma v\rangle=10^{-35}$ cm3/s.
Figure 4: Expected ionization rate in the CMZ from annihilation of sub-GeV DM particles into $e^{+}e^{-}$, as obtained with the numerical calculation for different DM distributions. The band illustrates the range of observed values of ionization rate in the CMZ.
Appendix C X and gamma ray emissions from annihilation of sub-GeV DM
In the left panel of Fig. 5, we compare the predicted X-ray emission (coming mostly from inverse Compton scattering of the DM-induced $e^{\pm}$ population with CMB and other interstellar radiation fields) from the annihilation of $1$ and $10$ MeV DM for different profiles and for $\langle\sigma v\rangle=10^{-35}$ cm3/s, with the most constraining data set of the XMM-Newton measurements derived in Ref. Foster et al. (2021), the Ring 3 De la Torre Luque et al. (2023d, c). The emission from the $e^{\pm}$ particles injected by $10$ MeV DM peak at a few keV, while the peak emission for the $1$ MeV DM products peaks below the keV Cirelli et al. (2023b); De la Torre Luque et al. (2023d). Since the infrared, optical and UV radiation fields are more concentrated around the GC, the inverse-Compton emission from the scattering with these fields is more intense for the cuspier DM profiles. As a result, annihilation of sub-GeV DM can explain the observed ionization rate at the CMZ without being in conflict with X-ray emissions.
Similarly, in the right panel of this figure we show the predicted MeV emission, dominated by bremsstrahlung, which is calculated with the Hermes code Dundovic et al. (2021), and compare it to SPI (INTEGRAL) measurements of the diffuse gamma-ray flux.
Figure 5: Left panel: X-ray emission associated to the $e^{\pm}$ population from DM compared to XMM-Newton data in the most constraining ring (Ring 3) Foster et al. (2021). Right panel: Bremsstrahlung emission associated with the $e^{\pm}$ population injected by DM compared to the diffuse MeV gamma-ray data from SPI (INTEGRAL) Bouchet et al. (2010); Siegert et al. (2016) in the $|b|<15^{\circ}$, $|l|<30^{\circ}$ region. The gamma-ray flux at energies above the DM mass is due to reacceleration, that boosts the $e^{\pm}$ population to higher energies De la Torre Luque et al. (2023d). In both cases, these estimations are made for a $1$ and $10$ MeV DM particle for $\langle\sigma v\rangle=10^{-32}$ cm3/s, for different DM profiles